Prove that for all integer n3, P(n+1,3)−P(n,3)=3P(n,2).


1
Expert's answer
2021-10-15T09:34:45-0400

We shall prove by mathematical induction.

For "n=3,"

"P(4,3)-P(3,3)=24-6=18=3(6)=3P(3,2)"

Thus it is true for "n=3"

Assume it is true for "n=k,k\geq3." Then

"P(k+1,3)-P(k,3)=3P(k,2)"

We shall show that it is true for "n=k+1,k\geq 3"

That is, P(k+2,3)-P(k+1,3)=3P(k+1,2)

"P(k+2,3)-P(k+1,3)=\frac{(k+2)!}{(k-1)!}-\frac{(k+1)!}{(k-2)!}=k(k+1)(k+2)-k(k-1)(k+1)\\
=3k(k+1)"

For the RHS

"3P(k+1,2)=3\left(\frac{(k+1)!}{(k-1)!}\right)=3k(k+1)"

Since LHS=RHS, then it is true for "n=k+1,k\geq 3"

Hence, "P(n+1,3)-P(n,3)=3P(n,2), n\geq 3"